# LiRA: Light-Robust Adversary for Model-based Reinforcement Learning in Real World

### v1

@@ 3.3 Integration with light robustness
To address these remaining issues, the light robustness (Fischetti and Monaci, 2009) is integrated with the maximization problem of the above lower bound. The light robustness establishes a tolerance for performance degradation due to disturbances and assigns that constraint to the optimization problem. This allows for a more intuitive setting since the “relative” tolerance can be defined, in comparison to, for example, the degree of regularization to the prior and the absolute predictive performance of the model.
@@
This constraint can be converted into the corresponding regularization term via Lagrangian method with an auto-tuned gain $\lambda$, as shown in the literature (Kobayashi, 2023). In summary, the proposed LiRA solves the following optimization problem to suppress conservativeness while making the model moderately robust.
@@ 4.1 Restricted normalizing flows (RNF)
In LiRA, the disturbance-aware/marginalized models are learned independently in order to compare their prediction performance. By using conditional normalizing flows (CNF) (Papamakarios et al., 2021), the disturbance-marginalized model can be regarded as the base distribution and then efficiently converted to the disturbance-aware model by invertible transformations $g$ conditioned on the disturbance.
@@
However, the excellent representation capability of CNF is more than sufficient to compensate for prediction errors in the base distribution. As a result, the two models might cause very different predictions. Ideally, this is inappropriate because the prediction bias should not be induced by the disturbance, while the prediction variance can be increased.
Therefore, in LiRA, the above requirement (i.e. no prediction bias) is explicitly incorporated into the model design. That is, the mean of the disturbance-marginalized model is carried over to that of the disturbance-aware model. RNF proposed in the literature (Kobayashi and Aotani, 2023) (see Fig. 2) makes it possible with the following restrictions:
@@ 4.2 Hindsight reparameterization gradient (HRG)
The adversary optimization is a challenge when solving the min-max problem of eq. (6). That is, since it is indirectly involved in the prediction performance of models, the likelihood-ratio gradient (well-known as REINFORCE (Williams, 1992)) is forced to be employed (Mohaghegh Dolatabadi et al., 2020). However, this approach is known to be inefficient and sometimes unstable.
As shown in the literature (Kingma and Welling, 2014; Haarnoja et al., 2018), a probability distribution can be efficiently optimized by a direct gradient of values sampled from it (in this paper, the disturbance $d$). Such a so-called reparameterization gradient needs the computational graph in $d$ to $\phi$ (the parameters for adversary), and by applying GRL (Ganin et al., 2016), the adversarial learning defined as the corresponding min-max problem can be converted to the minimization problem simply. However, in LiRA, this reparameterization gradient is only available by storing $d$ in a buffer while retaining its computational graph, or by learning in a streaming manner without the buffer. The former wastes a lot of memory, and the latter significantly degrades efficiency. In addition, it should be noticed that generating $d$ at each learning time without storing it in the buffer is inappropriate in LiRA because $s^{\prime}$ and $r$ are dependent on the disturbance at the time they experienced, which varies from the newly generated one.
To address this issue in implementing LiRA, a new trick, HRG, is proposed in this paper, as illustrated in Fig. 3. Specifically, by focusing on the fact that the disturbance is box-constrained as $d\in\mathcal{D}=[-d^{\mathrm{max}},d^{\mathrm{max}}]$, the probability distribution model for the adversary is designed with CNF, $g(d;s\phi)$, and a uniform distribution as its base $\mathcal{U}(-d^{\mathrm{max}},d^{\mathrm{max}})$. Note that CNF is a popular choice for adversarial learning due to its high expressive capability (Mohaghegh Dolatabadi et al., 2020). Anyway, $d$ sampled from this adversary is stored in the buffer as a number with no computational graph. When learning, if $\hat{d}$, which is numerically equivalent to $d$, can be resampled from the adversary with $\phi$ different from the time $d$ sampled, $\hat{d}$ can be utilized for learning as it is since it is with the computational graph to $\phi$ and equivalent to $d$ numerically.
@@ 4.3 Midrange-mean balancing (MMB)
As mentioned above, to increase the robustness of the model (and MPC using it), rare events with low prediction accuracy should be prioritized to some extent (Aotani and Kobayashi, 2024). For adjusting the degree of prioritization, however, the previous work employed a meta-optimization method (Aotani et al., 2021), which requires an extra computational cost. In addition, due to the adversarial learning nature of LiRA, it is expected that the low-quality data for the model and the adversary, which should be prioritized, might conflict with each other.
@@
In addition to this, MMB considers midrange, which is a metric to represent central tendency like mean and median (Dodge, 2003), as defined below.
@@ 5.1 Task
Numerical simulations are conducted to verify the robot behaviors learned with the proposed LiRA. Mujoco (Todorov et al., 2012) is used as the simulator, and a task is to maximize the forward speed of a worm-type robot (see Fig. 5). This robot locates 6–8 joints at equal intervals on its straight body: 4 controllable joints on the head side and 2–4 joints for disturbance on the tail side. Note that the detailed task information with state, action, reward design are described in B.
@@
During learning, disturbances are generated by the agent’s adversarial disturbance generator, but the following three types of disturbances are prepared for testing. Note that the prior in this task is given as Gaussian (a.k.a. nominal noise), $\mathcal{N}(0,\sigma_{0}=0.2/3)$. The first is the nominal noise and is the weakest. The second and third composes several Brownian noises $\epsilon^{\mathrm{Brown}}$ as below.
@@
where $K=2$ is the number of noises mixtured and $k=3$ for the second and $k=6$ for the third denote the scale factors from the nominal noise. $\sigma(\cdot)$ calculates the sample standard deviation along to pre-generated $\tilde{d}_{t}$ ($t=1,\ldots,T_{\mathrm{max}}$ with $T_{\mathrm{max}}$ the maximum time step for episode), not along to dimension of $d$. Under the above disturbances, each model is tested 30 trials, and the interquartile mean of returns on the respective trials (referring to the literature (Agarwal et al., 2021)) is used as a metric for each model. In total, 12 models are learned under the same conditions with the respective random seeds, and then their noise-robust statistics (specifically, the best and worst cases are excluded at first, the outliers detected by the 1.5 IQR rule are also excluded [^1] Due to the limited number of models, the interquartile mean, which excludes half of the data, was not employed., the remaining results are averaged with 95 % confidence interval) is finally used as a metric for that conditions.
The specific learning conditions are in D. During learning/inference, the robot is controlled with AccelMPPI (Kobayashi and Fukumoto, 2022) (its conditions are in E), but exploration noises are added to actions only during learning. Note that $\lambda$, which corresponds to the adversary level, is give to be a scalar in the following simulations in order to focus on its tuning process and relationship between it and the sensitivity to disturbances, as mentioned above.
@@ 5.2 Results
The test results are shown in Fig. 6. It can be found that Nominal without adversarial learning achieved high performance with small disturbance, but it was vulnerable to large ones. Full, which always learned adversarially to the maximum extent possible, maintained its performance independent of the disturbance intensity, but its basic performance was low and conservative. Compared to Nominal, Proposal (a.k.a. LiRA) was able to suppress the performance degradation caused by the disturbance intensity, was less conservative than Full, and succeeded in achieving a well-balanced and stable learning.
@@
In order to satisfy this constraint, how $\lambda$ was automatically tuned by LiRA is depicted in Fig. 8 with its statistics summarized in Table 2. Note again that the smaller $\lambda$ is, the stronger the disturbance intensity. The fewer the number of disturbance joints and the smaller the original disturbance influence, the smaller $\lambda$ was, giving priority to adversarial learning. Conversely, as the number of joints for external disturbances increased, $\lambda$ became larger, suggesting that the disturbance intensity was suppressed to prevent the performance degradation.
In addition, $\lambda$ was once large from the beginning to the middle of learning progress, and after a while, it became smaller and converged to each value suitable for the task. This result suggests the emergence of some kind of curriculum that once simplifies the task to easily predictable situations and then makes the task more difficult by gradually increasing the disturbance. Indeed, it can be said that the derivation of LiRA is partially similar to the self-paced learning methods (Kumar et al., 2010; Klink et al., 2020), and therefore, LiRA might lead to such an additional value. Note that the disturbance-aware model in LiRA can be considered as an automatic pacemaker for such a curriculum.
@@ 6.1 Task
The details of this task configuration and learning/MPC conditions are all described in C, D, and E, respectively. In summary, the adversary applies virtual horizontal forces to the robot body as like (Shi et al., 2024), and then, LiRA learns the model for optimizing the horizontal stride and stiffness. The nominal-level disturbance has little effect on the robot, not allowing it to learn how to respond to the disturbance at all, while the disturbance near its maximum intensity, which is adversarily given, can easily make the robot topple, causing a high risk of malfunction [^2] Indeed, the excessive disturbance given during debugging caused the robot to tip over and damage the frame.. That is, both Nominal and Full are not suitable for this task. In contrast, LiRA is expected to safely achieve the moderate robustness by controlling the disturbance intensity so as not to cause such undesired situations.
@@ 7.1 State dependency of adversary level
Figure 15: $\lambda$ for states experienced (embedded into two dimensions by UMAP (McInnes et al., 2018)): $\lambda$ was initialized around $0.5$; as LiRA progressed, $\lambda$ was polarized to zero or one, and gradually $\lambda\simeq 0$ increased; the cluster of states with $\lambda\simeq 1$ was still left at the end of learning probably as the states losing balance, which are difficult to predict future states.
As mentioned above, the inverse of the adversary level, $\lambda$, can be learned as a state-dependent function, as in the real-world demonstration above. In this case, not only the time variation of $\lambda$ but also the state-dependent differences should be confirmed. Therefore, $\lambda$ at each learning progress is illustrated in Fig. 15 for the states in the buffer finally collected during the above demonstration, which are embedded into two dimensions by UMAP (McInnes et al., 2018). At the start of learning, $\lambda$ was initialized to be near $0.5$ in all states. $\lambda$ was polarized to zero or one as learning progressed, and $\lambda\simeq 0$ increased gradually. In other words, the inexperienced states at 40min were with $\lambda\simeq 1$ at that time, and then, they were encountered by the enhanced disturbances, shifting $\lambda$ to zero by learning the way of predicting the future states from them. However, even at the final stage 80min, there was still a cluster of states with $\lambda\simeq 1$. This cluster probably represents states out of balance and cannot increase their adversary level because of low predictability. As a result, disturbances that would intentionally tip over might be suppressed, making it easier to ensure safety.
Thus, the state-dependent $\lambda$ basically functions as expected, but some issues remain. First, at 40min, the states with $\lambda\simeq 0$ and $\lambda\simeq 1$ were mixed, suggesting that a slight change in state might cause a sudden change in $\lambda$. Since such a sudden change in $\lambda$ induces instability in learning, it is desirable to smooth the function of $\lambda$, for example, as in the literature (Kobayashi, 2022). Furthermore, in the cluster that seems to be out of balance, some states with $\lambda\simeq 0$ at 40min were reverted to $\lambda\simeq 1$ at 80min. Since this is considered to be a failure in the estimation of $\lambda$, it would be better to update $\lambda$ more carefully, for example, by considering uncertainty in parameter estimation based on Bayesian theory (Wenzel et al., 2020). Finally, although the initial value of $\lambda$ was near $0.5$, it would be more natural to start with the larger $\lambda$.
@@ 7.2 Practical and theoretical time efficiency
The above issues regarding time efficiency are mainly due to the practical implementation, but apart from that, further theoretical contributions can be considered. First, the exploration by the RL agent in this implementation relies on Gaussian noise optimized by MPC, but it is known that such a random exploration is not efficient (Wilson et al., 2021). As suggested in the previous studies such as (Seo et al., 2021), the exploration efficiency could be improved by actively selecting the actions that can transition to states with high uncertainty from the vicinity of its likely optimal actions. The adversary may also be better to prioritize the disturbances that transition to states where its adversary level is uncertain in conjunction with making $\lambda$ stochastic suggested above. By such, the adversary could reduce the unnecessary attacks.

### v2

@@ 3.3 Integration with light robustness
To address these remaining issues, the light robustness [14] is integrated with the maximization problem of the above lower bound. The light robustness establishes a tolerance for performance degradation due to disturbances and assigns that constraint to the optimization problem. This allows for a more intuitive setting since the “relative” tolerance can be defined, in comparison to, for example, the degree of regularization to the prior and the absolute predictive performance of the model.
@@
This constraint can be converted into the corresponding regularization term via Lagrangian method with an auto-tuned gain $\lambda$, as shown in the literature [15]. In summary, the proposed LiRA solves the following optimization problem to suppress conservativeness while making the model moderately robust.
@@ 4.1 Restricted normalizing flows (RNF)
In LiRA, the disturbance-aware/marginalized models are learned independently in order to compare their prediction performance. By using conditional normalizing flows (CNF) [49], the disturbance-marginalized model can be regarded as the base distribution and then efficiently converted to the disturbance-aware model by invertible transformations $g$ conditioned on the disturbance.
@@
However, the excellent representation capability of CNF is more than sufficient to compensate for prediction errors in the base distribution. As a result, the two models might cause very different predictions. Ideally, this is inappropriate because the prediction bias should not be induced by the disturbance, while the prediction variance can be increased. It is possible to add a regularization term such that the two models match. However, such a soft constraint would be difficult to adjust and the degree of divergence could be oscillatory, which is undesirable since LiRA works on the basis of the difference in predictive performance between the two.
Therefore, in LiRA, the above requirement (i.e. no prediction bias) is explicitly incorporated into the model design. That is, the mean of the disturbance-marginalized model is carried over to that of the disturbance-aware model. RNF proposed in the literature [46] (see Fig. 3) makes it possible with the following restrictions:
@@ 4.2 Hindsight reparameterization gradient (HRG)
The adversary optimization is a challenge when solving the min-max problem of eq. (6). That is, since it is indirectly involved in the prediction performance of models, the likelihood-ratio gradient (well-known as REINFORCE [50]) is forced to be employed [51]. However, this approach is known to be inefficient and sometimes unstable.
As shown in the literature [13, 52], a probability distribution can be efficiently optimized by a direct gradient of values sampled from it (in this paper, the disturbance $d$). Such a so-called reparameterization gradient needs the computational graph in $d$ to $\phi$ (the parameters for adversary), and by applying GRL [47], the adversarial learning defined as the corresponding min-max problem can be converted to the minimization problem simply. However, in LiRA, this reparameterization gradient is only available by storing $d$ in a buffer while retaining its computational graph, or by learning in a streaming manner without the buffer. The former wastes a lot of memory, and the latter significantly degrades efficiency. In addition, it should be noticed that generating $d$ at each learning time without storing it in the buffer is inappropriate in LiRA because $s^{\prime}$ and $r$ are dependent on the disturbance at the time they experienced, which varies from the newly generated one.
To address this issue in implementing LiRA, a new trick, HRG, is proposed in this paper, as illustrated in Fig. 4. Specifically, by focusing on the fact that the disturbance is box-constrained as $d\in\mathcal{D}=[-d^{\mathrm{max}},d^{\mathrm{max}}]$, the probability distribution model for the adversary is designed with CNF, $g(d;s\phi)$, and a uniform distribution as its base $\mathcal{U}(-d^{\mathrm{max}},d^{\mathrm{max}})$. Note that CNF is a popular choice for adversarial learning due to its high expressive capability [51]. Anyway, $d$ sampled from this adversary is stored in the buffer as a number with no computational graph. When learning, if $\hat{d}$, which is numerically equivalent to $d$, can be resampled from the adversary with $\phi$ different from the time $d$ sampled, $\hat{d}$ can be utilized for learning as it is since it is with the computational graph to $\phi$ and equivalent to $d$ numerically.
@@ 4.3 Midrange-mean balancing (MMB)
As mentioned above, to increase the robustness of the model (and MPC using it), rare events with low prediction accuracy should be prioritized to some extent [45]. For adjusting the degree of prioritization, however, the previous work employed a meta-optimization method [53], which requires an extra computational cost. In addition, due to the adversarial learning nature of LiRA, it is expected that the low-quality data for the model and the adversary, which should be prioritized, might conflict with each other.
@@
In addition to this, MMB considers midrange, which is a metric to represent central tendency like mean and median [54], as defined below.
@@ 5.1 Task
Numerical simulations are conducted to verify the robot behaviors learned with the proposed LiRA. Mujoco [16] is used as the simulator, and a task is to maximize the forward speed of a worm-type robot (see Fig. 6). This robot locates 6–8 joints at equal intervals on its straight body: 4 controllable joints on the head side and 2–4 joints for disturbance on the tail side. Note that the detailed task information with state, action, reward design are described in B.
@@
During learning, disturbances are generated by the agent’s adversarial disturbance generator, but artificial disturbances must be applied to test the robustness of the learned policies. Although a realistic disturbance could be a complex terrain surface, for reproducibility, artificial noises are added to the joints for the disturbances. In this paper, the following three types of disturbances are prepared for testing. Note that the prior in this task is given as Gaussian (a.k.a. nominal noise), $\mathcal{N}(0,\sigma_{0}=0.2/3)$. The first is the nominal noise and is the weakest. It is confirmed that such Gaussian noise tends to cancel each other out in the time evolution even when the scale is increased, and it is difficult to become the disturbances that significantly change the tail’s posture. Therefore, the second and third composes several Brownian noises $\epsilon^{\mathrm{Brown}}$, which have strong low-frequency components effective in changing the tail’s posture, as below.
@@
where $K=2$ is the number of noises mixtured and $k=3$ for the second and $k=6$ for the third denote the scale factors from the nominal noise. $\sigma(\cdot)$ calculates the sample standard deviation along to pre-generated $\tilde{d}_{t}$ ($t=1,\ldots,T_{\mathrm{max}}$ with $T_{\mathrm{max}}$ the maximum time step for episode), not along to dimension of $d$. Under the above disturbances, each model is tested 30 trials, and the interquartile mean of returns on the respective trials (referring to the literature [55]) is used as a metric for each model. In total, 12 models are learned under the same conditions with the respective random seeds, and then their noise-robust statistics (specifically, the best and worst cases are excluded at first, the outliers detected by the 1.5 IQR rule are also excluded [^1] Due to the limited number of models, the interquartile mean, which excludes half of the data, was not employed., the remaining results are averaged with 95 % confidence interval) is finally used as a metric for that conditions.
The specific learning conditions are in D. During learning/inference, the robot is controlled with AccelMPPI [39] (its conditions are in E), but exploration noises are added to actions only during learning. Note that $\lambda$, which corresponds to the adversary level, is give to be a scalar in the following simulations in order to focus on its tuning process and relationship between it and the sensitivity to disturbances, as mentioned above.
@@ 5.2 Results
Figure 7: Learning curves of score (i.e. return): the implemented RL algorithm and adversarial learning work as expected.
@@
Table 2: Statistical performance evaluation by eq. (15)
Method Tail2 Tail3 Tail4
Nominal -1044$\pm$349 -768$\pm$265 -556$\pm$167
Full -994$\pm$391 -759$\pm$327 -394$\pm$197
Proposal -689$\pm$276 -553$\pm$232 -360$\pm$117
@@
Note that at the top priority in this paper is to confirm that the theoretically-derived LiRA behaves properly, comparisons are limited to that special cases and various existing methods are omitted.
First, to confirm that the RL algorithm in this study works properly, learning curves for the score are decpited in Fig. 7. The low score at the beginning of the learning process increased after a certain number of episodes, indicating that the policy optimization by RL worked as expected. The fact that the increase was low for Full and decreases after the increase for Proposal was due to the adversarial attacks.
Then, the test results are shown in Fig. 8. It can be found that Nominal without adversarial learning achieved high performance with small disturbance, but it was vulnerable to large ones. Full, which always learned adversarially to the maximum extent possible, maintained its performance independent of the disturbance intensity, but its basic performance was low and conservative. Compared to Nominal, Proposal (a.k.a. LiRA) was able to suppress the performance degradation caused by the disturbance intensity, was less conservative than Full, and succeeded in achieving a well-balanced and stable learning.
The following metric is defined to quantitatively evaluate the conservativeness reduction and the robustness improvement.
$(S^{\mathrm{nominal}}-\max_{\mathrm{methods,models}}S^{\mathrm{nominal}})$
$+$ $\sum_{\mathrm{i\in noises}}(S^{\mathrm{i}}-\max_{\mathrm{models}}S^{\mathrm{nominal}})$ (15)
where, $S$ denotes the return. The results of this statistical evaluation are summarized in Table 2. Nominal and Full did not have high values in any of the tasks due to low robustness or excessive conservativeness. In contrast, Proposal had the highest values in all tasks, and in addition, its standard deviations were smaller than the others.
@@
In order to satisfy this constraint, how $\lambda$ was automatically tuned by LiRA is depicted in Fig. 10 with its statistics summarized in Table 3. Note again that the smaller $\lambda$ is, the stronger the disturbance intensity. The overall trend is that the fewer the number of disturbance joints and the smaller the original disturbance influence, the smaller $\lambda$ was, giving priority to adversarial learning. Conversely, as the number of joints for external disturbances increased, $\lambda$ became larger, suggesting that the disturbance intensity was suppressed to prevent the performance degradation.
Looking at the early training period, we see that $\lambda$ was updated toward zero once in every trial. This is due to the fact that the initialized adversary was not subjected to a significant collapse of the models’ prediction accuracy, and the the disturbance-aware model was also not able to properly reflect the effects of disturbances in its predictions. After getting proper functions of them, $\lambda$ was once large from the beginning to the middle of learning progress, and after a while, it became smaller and converged to each value suitable for the task. This result suggests the emergence of some kind of curriculum that once simplifies the task to easily predictable situations and then makes the task more difficult by gradually increasing the disturbance. Indeed, it can be said that the derivation of LiRA is partially similar to the self-paced learning methods [17, 18], and therefore, LiRA might lead to such an additional value. That is, the disturbance-aware model in LiRA can be considered as an automatic pacemaker for such a curriculum. As another characteristic in the learning process, it can be found from the confidence intervals that when $\lambda$ started to converge statistically in the second half of the learning, the fluctuation of $\lambda$ was active, implying that perhaps adversarial learning reached an unstable equilibrium.
@@ 6.1 Task
The details of this task configuration and learning/MPC conditions are all described in C, D, and E, respectively. In summary, the adversary applies virtual horizontal forces to the robot body as like [56], and then, LiRA learns the model for optimizing the horizontal stride and stiffness. The nominal-level disturbance sets so as to have little effect on the robot, not allowing it to learn how to respond to the disturbance at all, while the disturbance near its maximum intensity, which is adversarily given, can easily make the robot topple, causing a high risk of malfunction [^2] Indeed, the excessive disturbance given during debugging caused the robot to tip over and damage the frame.. That is, both Nominal and Full are not suitable for this task. In contrast, LiRA is expected to safely achieve the moderate robustness by controlling the disturbance intensity so as not to cause such undesired situations.
@@ 7.1 State dependency of adversary level
Figure 17: $\lambda$ for states experienced (embedded into two dimensions by UMAP [57]): $\lambda$ was initialized around $0.5$; as LiRA progressed, $\lambda$ was polarized to zero or one, and gradually $\lambda\simeq 0$ increased; the cluster of states with $\lambda\simeq 1$ was still left at the end of learning probably as the states losing balance, which are difficult to predict future states.
As mentioned above, the inverse of the adversary level, $\lambda$, can be learned as a state-dependent function, as in the real-world demonstration above. In this case, not only the time variation of $\lambda$ but also the state-dependent differences should be confirmed. Therefore, $\lambda$ at each learning progress is illustrated in Fig. 17 for the states in the buffer finally collected during the above demonstration, which are embedded into two dimensions by UMAP [57]. At the start of learning, $\lambda$ was initialized to be near $0.5$ in all states. $\lambda$ was polarized to zero or one as learning progressed, and $\lambda\simeq 0$ increased gradually. In other words, the inexperienced states at 40min were with $\lambda\simeq 1$ at that time, and then, they were encountered by the enhanced disturbances, shifting $\lambda$ to zero by learning the way of predicting the future states from them. However, even at the final stage 80min, there was still a cluster of states with $\lambda\simeq 1$. This cluster probably represents states out of balance and cannot increase their adversary level because of low predictability. As a result, disturbances that would intentionally tip over might be suppressed, making it easier to ensure safety.
Thus, the state-dependent $\lambda$ basically functions as expected, but some issues remain. First, at 40min, the states with $\lambda\simeq 0$ and $\lambda\simeq 1$ were mixed, suggesting that a slight change in state might cause a sudden change in $\lambda$. Since such a sudden change in $\lambda$ induces instability in learning, it is desirable to smooth the function of $\lambda$, for example, as in the literature [58]. Furthermore, in the cluster that seems to be out of balance, some states with $\lambda\simeq 0$ at 40min were reverted to $\lambda\simeq 1$ at 80min. Since this is considered to be a failure in the estimation of $\lambda$, it would be better to update $\lambda$ more carefully, for example, by considering uncertainty in parameter estimation based on Bayesian theory [59]. Finally, although the initial value of $\lambda$ was near $0.5$, it would be more natural to start with the larger $\lambda$.
@@ 7.2 Practical and theoretical time efficiency
The above issues regarding time efficiency are mainly due to the practical implementation, but apart from that, further theoretical contributions can be considered. First, the exploration by the RL agent in this implementation relies on Gaussian noise optimized by MPC, but it is known that such a random exploration is not efficient [60]. As suggested in the previous studies such as [61], the exploration efficiency could be improved by actively selecting the actions that can transition to states with high uncertainty from the vicinity of its likely optimal actions. The adversary may also be better to prioritize the disturbances that transition to states where its adversary level is uncertain in conjunction with making $\lambda$ stochastic suggested above. By such, the adversary could reduce the unnecessary attacks.
7.3 Slowdown of auto-tuning in overlearning
In the real-robot experiment, LiRA succeeded in achieving the moderate robustness of the model by gradually increasing the disturbances. At the end of 80 minutes (i.e. 160 episodes), $\lambda$ had converged, but as suggested by the simulations, this is expected to be on an unstable equilibrium point. Therefore, it should be additionally checked whether this balance is lost when overlearning is performed, and whether that can be recovered by readjusting $\lambda$. Finally, 120 minutes (i.e. 240 episodes) of learning was conducted. As a result, the disturbances became more aggressive to make the model more unpredictable, causing the robot to lose its balance and fail episodes more frequently. Such a situation would normally be avoided by readjusting $\lambda$ so that the disturbances are again suppressed to restore balance. However, in this additional experiment, the increase in $\lambda$ became very slow and the disturbances could not be satisfactorily suppressed until the end.
The reason for this slowdown in the auto-tuning of $\lambda$ might be because the proportion of worst-case data in the buffer was smaller. In other words, the learning of normal data, which is the majority, inhibited the reflection of the worst-case data. Although MMB was expected to mitigate this situation, the worst cases were rarely included in small batches, and the worst cases could not be fully considered.
To address this limitation, a technique that more explicitly considers the value of the data is needed instead of (or in addition to) MMB. For example, several techniques have been proposed to prioritize the replay of data from the replay buffer [62, 63], and inspired by them, it would be effective to prioritize the replay of data with poor model prediction accuracy (or data with insufficient interference by the adversary). Alternatively, a weighting that is theoretically equivalent to the prioritization of replay data would also prevent the slowdown. One solution may be a method that utilizes the softmax operator as the weighting [64].
7.4 Heuristic design
In the theoretical derivation and practical implementation of LiRA, several hyperparameters need to be specified. This paper set them empirically.
In particular, the tolerance of performance degradation between the disturbance-aware and -marginalized models, $\rho$, has empirical lower and upper bounds, and then it was searched within that range for obtaining a reasonable $\lambda$ in the Tail3 simulation. In other words, the lower bound of $\rho$ is given to be the degradation when no disturbance is given, which is attributed to the additional expressive power of RNF. The upper bound of $\rho$ is roughly found based on the degradation when $\lambda$ is fixed to zero. In this way, the search range of $\rho$ can be limited, and the determined $\rho=1.5$ worked without problems even in other tasks. However, it is true that such a manual search is burdensome, although specifying $\rho$ is easier than restricting the disturbance intensity $d^{\mathrm{max}}$ and LiRA can adjust the adversarial level in a state-dependent manner. For example, the upper bound of $\rho$ could be made more user-specifiable by connecting it to acceptable “control” performance degradation, thus narrowing the search range beyond the current approach. It would be desirable to achieve this connection analytically, but even if the relationship is too complex, it would be possible by introducing meta optimization methods [65].
The other important hyperparameters are $\beta$ and $\epsilon$. $\beta$ determines the strength of the regularization of the adversary to its prior, and as $\lambda$ increases, this regularization becomes more dominant than adversarial learning, simplifying the adversary. To make LiRA a more general technology, it would be useful to conduct a deeper analysis of the adversary’s learning behavior according to $\lambda$ and $\beta$, and to design $\beta$ as a function of $\lambda$. On the other hand, $\epsilon$ is a threshold value for the slack variable to convert inequality into equality constraints for $\lambda$ optimization. When $\epsilon$ is small, the equality constraint is achieved through the slack variable optimization and automatic adjustment of $\lambda$ becomes slow. In this paper, $\epsilon$ is set relatively large so that the behavior of $\lambda$ can be easily observed, but the change speed of $\lambda$ is expected to have some effects on learning. For example, if it is too large, the learning process may become unstable because the optimization for a given $\lambda$ cannot be completed in time, or if it is too small, the sample efficiency may decrease due to the increase of similar data. Self-paced learning [66], in which $\epsilon$ is gradually increased to the extent that learning does not break down, may be one solution.
Although these hyperparameters have to be designed to be task-independent and generic, in real-world applications, the way of applying the disturbances and the generative model of the disturbances need to be carefully considered. The most effective way to do this is to reflect the domain knowledge of the target, but in other cases, interrupting the agent’s actions stochastically, as performed in the context of action robust RL [67], is a good and easy method to implement. In this case, the generative model can be designed on the same action space as the policy, and in the non-adversarial case (i.e. $\lambda\to 1$), the agent’s behavior should be passed directly to the environment. Such a behavior is available if the adversary’s prior is set to be the agent’s policy, or if the interrupting probability is also included in the generative model and its prior is designed to have the minimum probability. Note, however, that in the latter case, HRG is not applicable because the intervention signal is discrete. Anyway, to maximize the performance of LiRA, these heuristics still need to be designed for real-world applications.